# Learning Geospatial Region Embedding with Heterogeneous Graph

### v1

@@ 2 Related Work
Geospatial Embedding. Numerous prior studies [52, 57, 13, 35, 20] have attempted to address the challenge of geospatial embedding. For example, Zhang et al. [52] proposed a multi-view graph representation approach that considered POI data and mobility patterns to generate representations of regions. In a similar vein, Zhou et al. [57] employed a prompt learning method by leveraging both POI and mobility data. However, these approaches are limited to specific tasks and regions due to limited and non-comprehensive embeddings [53, 42]. To tackle this issue, there has been a growing interest in utilizing satellite imagery for geospatial embedding, as it offers easy accessibility and global coverage. For instance, PG-SimCLR [38] and UrbanCLIP [42] have demonstrated success in profiling urban regions through satellite images for various downstream tasks. Nevertheless, there remains a gap for these models to effectively interpret the complex and professional semantic meaning of satellite imagery. Moreover, the aforementioned methods overlook the crucial modeling of mixed-order inter-region relations, which could affect the comprehensiveness of geospatial embedding. Therefore, our proposed GeoHG initially employs a heterogeneous graph to capture these inter-region relationships while concurrently learning the explicit features of a given region.
Graph Neural Network for Geospatial Representation. Graph Neural Networks (GNNs) offer a succinct and scalable approach for modeling intricate geospaces with non-Euclidean characteristics [51, 26, 11, 34]. By propagating messages through edges, GNNs are capable of learning the geospatial correlations between regions and geo-entities effectively. To better represent these complex correlations, numerous studies have adopted a multi-view graph structure [22, 5, 3, 7]. In this framework, nodes in each view (e.g., distance, mobility, semantic) are linked with distinct edge vectors that represent the correlations between node pairs under specific perspectives. Differing from this approach, our work utilizes a Heterogeneous Graph (HG) to further represent mixed-order geospatial correlations. Although HG structures are frequently used in knowledge graphs [27, 17, 23] that often lack scalability and effective training [27, 17], our GeoHG innovatively revises HG to succinctly represent mixed-order correlations and group interactions between regions from both environmental and societal perspectives. Moreover, we devise a streamlined yet efficient pipeline for automatically constructing this heterogeneous graph from satellite imagery and POI information.
Multimodal Learning in GeoAI. Multimodal learning involves integrating data from various modalities to enhance model performance [1, 8, 6]. Geospatial multimodal learning, in particular, improves regional understanding by combining spatio-temporal (e.g., POI and road network), visual (e.g., urban imagery), textual (e.g., social media data) modalities, resulting in more comprehensive and accurate geospatial representations [59, 55]. Traditional geospatial representation learning mostly relies on task-specific supervised methods [48, 55], which are limited by their dependence on domain expertise and inability to adapt to new tasks [59]. To overcome these limitations, recent studies [15, 52] focus on learning general geospatial embeddings from a multimodal perspective. Nonetheless, efficiently integrating multimodal data from intricate geospaces, while maintaining generalizability, poses a significant challenge [59]. In this study, we develop a holistic geospatial embedding learning framework grounded in graph theory to amalgamate visual insights and POI information of geospace with geospatial mixed-order relation awareness while maintaining awareness of geospatial mixed-order relations.
3 Methodology
Geospatial Region Embedding: We aim to train a GeoHG model capable of generating embeddings for regions worldwide to facilitate a diverse set of downstream applications. To enable our proposed model to derive embeddings for arbitrary regions, we construct the GeoHG architecture using globally available satellite imagery and readily accessible POI data as input sources, serving as proxies for environmental and societal information, respectively.
Given the access to satellite imagery and POI dataset $D_{\rm satellite\_POI}=\left\{\left(S_{i},P_{i}\right)\right\}_{i=1}^{I}$ containing pairs of satellite image $S_{i}$ and POI data $P_{i}$ corresponding to the region of interest $R_{i}$, our GeoHG model W generates embedding $E_{i}=\phi\left(S_{i},P_{i}\right)$, where $E_{i}\in\mathbb{R}^{d}$, representing the respective region embedding. $\phi(\cdot)$ is an embedding function capturing semantic and contextual information of $R_{i}$. The generated embeddings $E_{i}$ encapsulate the intra-region features containing environmental and societal information and its higher-order inter-region dependency with other regions. Notably, the satellite images $S_{i}$ and $P_{i}$ can originate from arbitrary geographical regions worldwide, necessitating that the learned embeddings concisely yet comprehensively encode rich contextual information across a global scale, which poses a formidable challenge.
Figure 2: GeoHG Framework. (a) Given satellite images and POI information, we derive intra-region features based on geo-entities proportion, POI categories proportion, and location. (b) To model inter-region correlations, we construct a heterogeneous graph structure that reflects the mixed-order relationships from regions’ spatial, environmental and societal perspectives. Details are displayed in Appendix C.2. (c) We integrate the intra-region features and inter-region correlations through a model-agnostic graph architecture. (d) We propose two training strategies: self-supervised pretraining and end-to-end training on specific tasks.
Our architecture comprises four major stages to obtain intra-regional and inter-regional correlations and effectively integrate them, as illustrated in Figure 2: intra-region feature representation, inter-region feature representation, heterogenous graph-based representation integration and pre-training and end-to-end training. Now, we present the details of each stage.
3.1 Intra-Region Feature Representation
Figure 3: Spatial Grids.
To effectively encode spatial, environmental, and societal signals into intra-region representations, we extract multi-modal geo-contextual data including satellite imagery, POI information, and the location of the region of interest.
Spatial Position Embedding: We adopt a simple yet effective method to extract the spatial information of the region of interest, drawing inspiration from industrial conventions [42, 9]. Specifically, we divide the overall geospace into multiple $1\textrm{km}\times 1\textrm{km}$ spatial grids with coordinate system origin ($lon^{0},lat^{0}$) and use the abscissa $x_{R}$ and ordinate $y_{R}$ of the region of interest $R$ as spatial information $E_{pos}(x,y)$, illustrated in Figure 3. Notably, $E_{pos}$ is directly related to the geo-coordinate ($lon,lat$) at the centroid of the targeted region, for any region $E^{i}$ in the target space, we have:
$E^{i}_{pos}=(lon^{i}-lon^{0},lat^{i}-lat^{0})\Delta*D^{-1}$ (1)
where $\Delta$ is the transform vector decided by the coordinate system. $D$ denotes the scale of the grids, and for $1\textrm{km}\times 1\textrm{km}$ grids, $D=1$.
Environmental Feature Embedding: We utilize satellite imagery to mine the environmental features and expect the satellite imagery encoding to be efficient and concise. To achieve this, we draw inspiration from the European Space Agency (ESA) with ESA WorldCover Dataset [50], elaborate in Appendix A, and design a segmentation-based process to encode the satellite images instead of the most commonly used visual encoders such as CNNs. Given a satellite image $S_{i}$ of a region of interest $R_{i}$, we conduct semantic segmentation to obtain a series of geo-entities $\{entity_{1},entity_{2},...,entity_{j}\}_{j=1}^{J}$ within that region, such as developed areas, grass, and trees based on ESA Worldcover, leveraging their area proportion as environmental feature embedding as below:
$E_{env}=\{p_{ent_{1}},p_{ent_{2}},...,p_{ent_{j}}\}_{j=1}^{J}$ (2)
where $p_{ent_{j}}=\frac{A_{j}}{A_{R}}$, and $A_{j}$ is the area occupied by $entity_{j}$, $A_{R}$ is the area of region $R_{i}$, $J$ is the number of environmental entity types. We discuss the motivation for using the segmentation-based approach instead of visual encoders and additional efficiency experiments in Appendix B.
$E_{soc}^{\prime}=\{p_{poi_{1}},p_{poi_{2}},...,p_{poi_{k}}\}_{k=1}^{K},\quad p_{\text{poi}_{k}}=|C_{k}\cap D_{POI}^{R_{i}}|/|D_{POI}^{R_{i}}|$ (3)
where $K$ is the number of POI categories, $C_{k}=\{poi_{j}\in D_{POI}\mid\text{category}(poi_{j})=k\}$. However, directly employing the proportion of different POI categories $E_{soc}^{\prime}$ as societal features can still be problematic, since it cannot distinguish between areas with high POI density (usually with a higher degree of socialization) and those with relatively low POI density. Therefore, we transform $E_{soc}^{\prime}$ into final societal feature embedding $E_{soc}$ by multiply a social impact factor $f$ as below:
$E_{soc}=f\cdot E_{soc}^{\prime},\quad f=log(D^{R_{i}}_{POI}+1)$ (4)
Finally, the intra-region feature embedding is generated by concatenating the spatial position embedding $E_{pos}$, environmental feature embedding $E_{env}$ and societal feature embedding $E_{soc}$:
$E_{intra}=concat[E_{pos},E_{env},E_{soc}]$ (5)
3.2 Inter-Region Correlation Representation
Our objective is to model pairwise second-order relationships between regions and their adjacent neighbors and higher-order dependencies with distant region groups exhibiting similar environmental or societal characteristics. Inspired by hypergraph theories [21, 4], elaborated in Appendix C.1, we employ a heterogeneous graph formulation GeoHG to represent these relationships explicitly. As shown in Figure 4, we take heterogeneous nodes in the graph structure as transfer nodes to achieve high-order message passing in the hypergraph. Given a pair of $1\textrm{km}\times 1\textrm{km}$ satellite image $S_{i}$ and POI data $P_{i}$ in $D_{satellite\_{POI}}$, we construct an undirected weighted heterogeneous graph $G_{i}$, containing three types of nodes: regional nodes, environmental entity nodes, and societal entity nodes. Specifically, regional nodes correspond to $1\textrm{km}\times 1\textrm{km}$ grid cells from the geospace; environmental entity nodes represent the $J$ distinct geo-entity classes detected from satellite image $S_{i}$, which are consistent with intra-region environmental features in Section 3.1; societal entity nodes constitute the $K$ different POI categories. The overall structure of GeoHG is shown in Figure 2 and the details are discussed in Appendix 4. We then elaborate on how second-order and high-order information is obtained through GeoHG $\mathcal{G}_{i}=(\mathcal{V},\mathcal{E})$, respectively.
Second-Order Relation Representation: Second-order relations capture the pairwise spatial adjacency between regional nodes (grid cells) in $\mathcal{G}_{i}$. We construct undirected edges $\mathcal{E}_{RNR}$, Region Nearby Region (RNR) between regional nodes whose corresponding grid cells are spatially adjacent in a $3\times 3$ grid, thereby encoding these local second-order dependencies.
High-Order Relation Representation: Differ from traditional multi-view graphs [7, 59] represent complex correlations through additional second-order region connections, we explicitly model higher-order relations by leveraging the environmental entity nodes $\mathcal{V}_{env}$ and societal entity nodes $\mathcal{V}_{soc}$ in $\mathcal{G}_{i}$. Specifically, we construct heterogeneous weighted edges $\mathcal{E}_{ELR}$ and $\mathcal{E}_{SLR}$ through $\mathcal{V}_{env}$ and $\mathcal{V}_{soc}$ to associate each regional node with its constituting environmental entities (e.g., water, vegetation) and societal entities (e.g., educational, commercial POIs). For $p_{ent_{it}}\geq\theta_{Env}$ or $f\cdot p_{poi_{ik}}\geq\theta_{Soc}$:
where $p_{ent_{ij}}$ represents the area proportion of geo-entity $j$ at region $i$, $p_{poi_{ij}}$ denotes the quantity proportion of POI category $k$ at region $i$ and $f$ is the social impact factor elaborated in Equation 4. $\theta_{Env}$ and $\theta_{Soc}$ serve as hyperparameters designed to optimize the graph structure for conciseness and efficiency. This edge formulation allows encoding higher-order relationships between regions exhibiting similar environmental/societal attributes, even if they lack spatial proximity. By combining the second-order spatial adjacency edges with these higher-order hyperedge associations, $\mathcal{G}_{i}$ holistically represents the mixed-order relational patterns within the region.
3.3 Heterogenous Graph-Based Representation Integration
Having derived intra-region features capturing regions’ intrinsic attributes and inter-regional dependencies encoded in the heterogeneous graph $\mathcal{G}_{i}$, we employ a model-agnostic graph neural network framework to jointly reason over the intra-regional and inter-regional information. In the graph, each regional node $v$ in $\mathcal{G}_{i}$ is represented with an initial node feature $\mathbf{x}_{v}$ corresponding to its intra-regional feature $E_{intra}$ elaborated in Sec 3.1. Note that for simplicity, instead of utilizing hypergraph neural networks [4, 44, 39], we model hyperedges with the form of complete subgraphs in $\mathcal{G}_{i}$. Therefore, we apply Heterogeneous Graph Neural Networks (HGNNs) over $\mathcal{G}_{i}$ to update the node representations by iteratively aggregating and transforming multi-hop neighborhood information:
where $\mathcal{N}(v)$ denotes the neighbors (spatial and higher-order) of node $v$ in $\mathcal{G}_{i}$. After L layers of updates, the final node embedding $\mathbf{x}_{v}^{(L)}$ integrates cues from the intra-regional features as well as mixed-order inter-region dependencies. This final embedding $\mathbf{x}_{v}^{(L)}$ serves as the comprehensive representation integrating intra-regional and inter-regional information for the given region node. Our framework is model-agnostic, enabling flexible integration of any HGNN variant for neighborhood aggregation, and allowing easy extension to more expressive HGNN models.
3.4 Pretraining and End-to-End Training
To effectively empower various geospatial tasks with the devised representations, we present two training strategies: self-supervised pre-training that learns generalizable region representations without task-specific labels, enabling efficient transfer to diverse downstream tasks; and end-to-end training that directly optimizes task-specific objectives for peak performance.
Self-Supervised Pre-training: Inspired by CLIP [29], we pre-train our model using a contrastive learning paradigm where we maximize the similarity between a region $\mathbf{R}_{i}$ and its corresponding positive region groups $\mathbf{C}_{i}$, while minimizing its similarity with other region groups in the batch. Specifically, we sample regional nodes adjacent to $\mathbf{R}_{i}$ as well as nodes exhibiting similar intra-regional features to construct $\mathbf{C}_{i}$. During training, we leverage a GNN model $\mathcal{M}_{pretrain}$ to obtain embeddings for $\mathbf{R}_{i}$ and all nodes in $\mathbf{C}_{i}$. The embeddings of the sampled positive regional nodes are further pooled to obtain the unified positive region representation $\mathbf{e}_{j}$. The process of generating embeddings for $\mathbf{R}_{i}$ and $\mathbf{C}_{i}$ is conducted as follows:
We further define a similarity function $f_{\mathbf{score}}$ to measure the similarity between the representations $\mathbf{e}_{i},\mathbf{e}_{j}$. $f_{\mathbf{score}}$ can be a simple dot product or a more complex metric and then employ an InfoNCE-based loss [28] to conduct contrastive learning:
where $f_{\mathbf{score}}(\mathbf{e}_{i},\mathbf{e}_{j})$ represents the score of positive pairs while $f_{\mathbf{score}}\left(\mathbf{e}_{i},\mathbf{e}_{n}\right)$ refers to the scores of negative pairs. After obtaining the pre-trained model $\mathcal{M}_{pretrain}$ through self-supervised learning, we fine-tune it for different downstream tasks by adding three trainable linear layers. The weights of these additional linear layers are updated during training for the specific downstream task while the weights of $\mathcal{M}_{pretrain}$ itself remain fixed.
End-to-End Training: For a specific downstream task, we directly optimize the whole GeoHG’s parameters from scratch using the task’s supervised signal. Taking the regional regression task as an illustrative example, we adopt the HGNN model $\mathcal{M}$ to obtain embeddings for given region $i$, then feed it into a three-linear layer regression head. $\mathcal{M}$’s parameters are updated by minimizing the Mean Square Error (MSE) loss on the training data, jointly learning task-specific region representations.
4 Experiments
Datasets & Baselines. The datasets used in this paper include satellite imagery, POI information and five region indicators (Population, GDP, Night Light, Carbon, $PM_{2.5}$) located in four representative cities in China: Beijing, Shanghai, Guangzhou and Shenzhen. The satellite images are collected according to geospatial grids and each presents a spatial area of 1 $km^{2}$. The entity segmentation results for satellite imagery are collected from the European Space Agency. We randomly split the data into 60% training, 20% validation, and 20% testing sets. For comparison, we select two classical methods (AutoEncoder [18] and ResNet-18 [10]), the state-of-the-art satellite imagery-based models ( UrbanCLIP [42]) and multisource and multimodal approaches (UrbanVLP [9], PG-SimCLR [38] and GeoStructual [20]) for geospatial region embedding. Moreover, for optimal generalization capability, we explore a variation of GeoHG where we employ contrast-based Graph Self-Supervised Learning (SSL) pretraining instead of end-to-end training, represented as GeoHG-SSL. Dataset, baseline and implementation details are provided in Appendix D, E.
4.1 Comparison with State-of-the-Art Methods
Table 1: Region indicators prediction results. The bold/underlined font means the best/the second-best result.
Methods GeoHG GeoHG-SSL UrbanVLP GeoStructural PG-SimCLR UrbanCLIP ResNet-18 AutoEncoder
Metric $\text{R}^{2}$ MAE $\text{R}^{2}$ MAE $\text{R}^{2}$ MAE $\text{R}^{2}$ MAE $\text{R}^{2}$ MAE $\text{R}^{2}$ MAE $\text{R}^{2}$ MAE $\text{R}^{2}$ MAE
Beijing Carbon 0.954 0.110 0.937 0.161 0.787 0.353 0.765 0.378 0.442 0.631 0.664 0.528 0.394 0.577 0.298 0.565
Population 0.874 0.271 0.870 0.282 0.725 0.404 0.730 0.402 0.471 0.964 0.461 0.552 0.266 0.623 0.168 0.667
GDP 0.647 0.336 0.644 0.331 0.586 0.416 0.617 0.401 0.277 0.768 0.355 0.539 0.285 0.699 0.171 0.822
Night Light 0.901 0.239 0.900 0.244 0.531 0.394 0.488 0.429 0.369 0.404 0.420 0.457 0.348 0.576 0.276 0.643
$PM_{2.5}$ 0.971 0.064 0.970 0.065 0.641 0.484 0.694 0.306 0.398 0.624 0.533 0.556 0.341 0.589 0.209 0.659
Shanghai Carbon 0.915 0.157 0.912 0.162 0.716 0.392 0.688 0.413 0.298 0.712 0.671 0.426 0.326 0.465 0.230 0.532
Population 0.936 0.161 0.928 0.172 0.593 0.471 0.613 0.456 0.315 0.731 0.456 0.557 0.279 0.627 0.166 0.742
GDP 0.778 0.323 0.767 0.331 0.310 0.595 0.377 0.553 0.294 0.767 0.326 0.587 0.289 0.711 0.197 0.844
Night Light 0.898 0.222 0.891 0.234 0.457 0.494 0.442 0.517 0.308 0.566 0.387 0.511 0.244 0.571 0.164 0.617
$PM_{2.5}$ 0.866 0.120 0.836 0.150 0.486 0.497 0.527 0.398 0.303 0.617 0.444 0.518 0.292 0.642 0.243 0.691
Guangzhou Carbon 0.885 0.219 0.884 0.209 0.698 0.385 0.681 0.497 0.422 0.708 0.585 0.444 0.375 0.515 0.254 0.570
Population 0.871 0.244 0.855 0.255 0.665 0.441 0.687 0.433 0.303 0.954 0.533 0.567 0.274 0.671 0.195 0.753
GDP 0.715 0.371 0.712 0.366 0.436 0.541 0.439 0.533 0.282 0.897 0.440 0.546 0.251 0.725 0.176 0.811
Night Light 0.871 0.234 0.854 0.248 0.577 0.418 0.574 0.415 0.435 0.433 0.483 0.478 0.242 0.551 0.176 0.602
$PM_{2.5}$ 0.833 0.133 0.822 0.158 0.638 0.462 0.652 0.461 0.315 0.542 0.56 0.514 0.231 0.694 0.196 0.780
Shenzhen Carbon 0.926 0.128 0.912 0.162 0.659 0.418 0.647 0.431 0.257 0.683 0.562 0.483 0.241 0.577 0.189 0.634
Population 0.892 0.165 0.879 0.173 0.790 0.343 0.797 0.314 0.311 0.758 0.527 0.592 0.299 0.654 0.175 0.772
GDP 0.798 0.297 0.767 0.331 0.532 0.448 0.517 0.455 0.307 0.895 0.508 0.464 0.234 0.817 0.119 0.884
Night Light 0.942 0.149 0.939 0.155 0.457 0.459 0.445 0.488 0.454 0.358 0.387 0.511 0.243 0.543 0.166 0.608
$PM_{2.5}$ 0.906 0.116 0.905 0.117 0.566 0.494 0.597 0.451 0.323 0.613 0.430 0.586 0.273 0.598 0.149 0.645
We adopt Mean Absolute Error (MAE), rooted mean squared error (RMSE), and coefficient of determination ($R^{2}$) as evaluation metrics. An increase in $R^{2}$, along with a reduction in MAE and RMSE values is indicative of enhanced model accuracy. We report the performances of each model in Table 1 and the table with RMSE metrics is provided in Appendix F. From these tables, we have three key findings: 1) Both GeoHG and GeoHG-SSL outperform all competing baselines over the 5 datasets for 4 cities. For instance, GeoHG surpasses the previous SOTA performance in Beijing, achieving $R_{2}$ improvements of +16.7%, +14.9%, +6.1%, +37%, +33% in Carbon, Population, GDP, Night Light and $PM_{2.5}$ tasks respectively. Consistent trends are observed across other cities, underscoring GeoHG’s stable accuracy, versatility, and strong generalization capabilities for geospatial region embedding. 2) The end-to-end trained GeoHG outperforms its pre-trained version GeoHG-SSL in most tasks while GeoHG-SSL shows a good performance in multi-task transferring. 3) Multisource and multimodal approaches, i.e., UrbanVLP [9] and GeoStructual [20], largely surpass traditional satellite imagery-based models by their more comprehensive embedding views.
4.2 Ablation Study & Interpretation Analysis
Effects of Core Components. To examine the effectiveness of each core component in our proposed framework, we conducted an ablation study based on the following variants for comparison: a) w/o Env, which excludes environment information from the satellite imagery for embedding. b) w/o Soc, which omits POI entities for embedding. c) w/o Pos, which does not utilize the corresponding location information of the regions. The $R^{2}$ results for five tasks in two cities, Guangzhou and Shanghai, are displayed in Figure 5. We can observe that removing location information markedly degrades performance across all tasks. Meanwhile, environmental features severely impact model accuracy on the PM2.5, Night Light, and GDP tasks. In contrast, excluding POI data slightly reduces performance on the Carbon task. Appendix F.3 shows ablation results for the other two cities.
Figure 5: Results of Ablation Study on $R^{2}$ Metric.
Effect of Graph Construction. Our framework utilizes heterogeneous graph and hyperedges between regions to reflect the geospatial high-order relations. To validate the effectiveness of high-order relation representation, we compare GeoHG against two variants: GeoHG-MLP which discards graph structure and relies only on intra-region feature representation, employing a 3-layer MLP for regression; GeoHG-Mono which keeps edges in adjacent regions while discarding hyperedges for high-order relations. The results presented in Figure 5 indicated that our GeoHG significantly benefits from efficient representation of complex mixed-order relations in geospace and integration of intra-region and inter-region representations, thereby resulting in enhanced performance.
Figure 6: Visualization of Geospatial Correlations learned by GeoHG in Appendix F.5. Index $(x,y)$ of a region represents the numbering in the latitude and longitude directions.
Interpretation of Mixed-order Geospatial Correlations. To investigate the power of heterogeneous graph structure in capturing mixed-order geospatial relations, inspired by GNNExplainer [46], we depict the trend of learned geospatial dependency among Carbon Emission datasets in Guangzhou. Notably, we do not incorporate any external features (e.g., road network, human mobility data) to construct aided edges in the graph. By selecting the top $N$ important nodes by their weight in the GNN prediction process, as illustrated in Figure 6, we observe that GeoHG effectively captures the high-order correlations between the target region and remote region groups through message passing within the heterogeneous graph structure. The details about GNNExplainer and experiment results are introduced in Appendix F.5.
4.3 Analysis on Few-shot Learning and Data Efficiency
Applying geospatial models globally across tasks often requires large labeled datasets for supervised training - a process that is time-consuming, computationally expensive, and hindered by data scarcity. We therefore evaluate GeoHG’s performance under limited data regimes of 5%, 10%, and 20% available training samples, with results shown in Figure 7. Results demonstrate GeoHG’s strong data efficiency - with only 5% data, GeoHG outperforms previous SOTA methods like UrbanVLP [9] and StructuralGeo [20] trained on the entire training dataset across all tasks. Utilizing only 20% of the data, GeoHG suffers minor performance degradation of 4%, 0.5%, 2.01%, 1.7%, and 3.2% on the Carbon, GDP, Light, PM2.5, and Population tasks, respectively. These findings illustrate GeoHG’s promising potential for data-efficient global deployment across diverse geospatial prediction tasks.
Figure 7: Left: Data Efficiency of GeoHG in $5$ region indicator prediction tasks. We take the average of $R^{2}$ results in 4 cities. Right: Visualization of Data Efficiency Result ($5\%$) in Population Prediction.
Geospatial data interpolation methods like Inverse Distance Weighting (IDW) [58, 24] and Universal Kriging (UK) are employed to generate predictions for data-scarce regions. We evaluate GeoHG against these methods on population prediction in an enlarged area of Shenzhen, with only 5% visible data (863 points), tasked with inferring the remaining 16,401 regions (16,401 $km^{2}$), as illustrated in Figure 7. GeoHG accurately captures the true distribution, substantially outperforming IDW and UK which deviate significantly. This highlights that traditional methods, solely relying on spatial relationships for interpolation, fail to model the intricate socio-environmental dependencies critical for characterizing population distributions. In contrast, GeoHG’s effective geospatial representations and modeling of high-order relationships enable accurate data pattern learning from limited samples.
5 Conclusion
In this paper, we propose GeoHG, a novel heterogeneous graph structure coupled with an efficient learning framework, specifically designed to generate informative geospatial embeddings for global regions. Our approach effectively captures comprehensive intra-region features from environmental and societal perspectives, as well as higher-order inter-region correlation through a heterogeneous graph formulation, and offers a seamless integration of these components within a model-agnostic graph structure. Extensive experiments across multiple datasets demonstrate GeoHG’s superior performance compared to existing methods. Notably, our method exhibits competitive performance even when the training data is significantly reduced. Due to the page limit, we provide more discussion in Appendix H, including the limitations, future directions and the social impact of our research.
Supplementary for: “GeoHG: Learning Geospatial Region Embedding with Heterogeneous Graph”
We organize our supplementary document as follows:
  1. A
    More Introduction of ESA WorldCover Dataset
  2. B
    Motivation for using Entity Segmentation for Satellite Imagery Encoding
    1. 1
      Introduction of semantic segmentation-based approach
    2. 2
      Validation experiment for segmentation-based approach
  3. C
    More Details of GeoHG
    1. 1
      Basic graph theory and our motivation
    2. 2
      Details of GeoHG structure
  4. D
    Dataset and Experiment Settings
  5. E
    Details of Baselines
  6. F
    More Details about Experiments Results
    1. 1
      RMSE metrics of experiment results
    2. 2
      Mean and standard deviation of metrics for GeoHG
    3. 3
      More ablation study results
    4. 4
      Details of data efficiency experiment
    5. 5
      Details of investigation of GeoHG with GNNExplainer
  7. G
    Qualitative Demonstration
  8. H
    More Discussion
Appendix A More Introduction of ESA WorldCover Dataset
The ESA WorldCover dataset  [^1] https://esa-worldcover.org/en represents a groundbreaking advancement in land use/land cover mapping, offering freely accessible, high-resolution (10 m) global coverage based on satellite imagery. Inspired by the 2017 WorldCover conference  [^2] https://worldcover2017.esa.int/, the European Space Agency (ESA) launched the WorldCover project and the primary accomplishment of this endeavor was the introduction in October 2021 of a freely accessible global WorldCover Dataset at a groundbreaking 10 m resolution for the year 2020 [49, 50]. This dataset leverages satellite imagery from both Sentinel-1 and Sentinel-2, and encompasses 11 distinct geo-entity categories, shown in Figure 8. It has also undergone rigorous independent validation by Wageningen University (for statistical accuracy) and the International Institute for Applied Systems Analysis (IIASA) (for spatial accuracy), attaining a notable global overall accuracy of approximately 75% [32].

### v2

@@ 5 Related Work
Deep Learning for Geospace Embedding. Numerous prior studies (Zhang et al., 2021; Zhou et al., 2023; Huang et al., 2023; Wang et al., 2020; Li et al., 2022) have tackled spatial embedding in data-driven solutions. For example, Zhang et al. (2021) used a multi-view graph representation that integrates POI data and mobility patterns, while Zhou et al. (2023) applied a prompt learning technique combining both POI and mobility data. Recent models like PG-SimCLR (Xi et al., 2022) and UrbanCLIP (Yan et al., 2024) have succeeded in profiling urban regions through satellite images. However, these models still struggle with indicator inference from limited samples and do not sufficiently explore the joint representation of geospace from geometric, environmental, and societal perspectives.
Socioeconomic Indicator Inference. Socioeconomic indicator inference is based on spatial interpolation methods, including Inverse Distance Weighting (IDW) (Zimmerman et al., 1999; Lu and Wong, 2008) and Kriging (Brus and Heuvelink, 2007), which are commonly used for this purpose. IDW is a straightforward method that weights sample points inversely by distance, but it does not consider spatial correlation and assumes a monotonous decrease in influence with distance (Mitas and Mitasova, 1999). In contrast, Kriging is a more complex method that incorporates both distance and spatial correlation, utilizing variograms and covariance functions to minimize prediction errors and provide uncertainty estimates. However, Kriging is computationally intensive for large spatial domains and relies on assumptions of spatial stationarity and isotropy (Leal-Villaseca et al., 2025). These widely-used interpolation methods are built on the assumptions of spatial continuity and isotropy.
6 Conclusion and Future Work
In this paper, we introduce GeoHG, a space-aware method for socioeconomic indicator inference that employs a heterogeneous graph-based geospace representation framework. Our approach enhances indicator inference from environmental and societal perspectives, as well as higher-order cross-region correlations. Extensive experiments demonstrate that GeoHG outperforms existing methods while maintaining strong performance even with limited data. Its minimal data reliance and double deployment strategies facilitate easy implementation for various domains. In the future, we plan to integrate advanced graph geometric techniques to accommodate more complex spatial structures within our graph framework, such as spherical space representation. This enhancement will allow the method to handle larger or global-scale inference where planar geometry is inadequate. Additionally, we aim to refine our code into a plug-and-play toolkit for indicator inference and interpolation, benefiting researchers in various fields.
Supplementary for:
“Space-aware Socioeconomic Indicator Inference with Heterogeneous Graphs”
Appendix A More Related Work
Multimodal Learning in GeoAI. Multimodal learning involves integrating data from various modalities to enhance model performance in various tasks [Baltrušaitis et al., 2018; Guo et al., 2019; Gao et al., 2020]. Traditional geospatial embedding learning mostly relies on task-specific supervised methods [Yuan et al., 2012; Zheng, 2015], which are limited by their dependence on domain expertise and inability to adapt to new tasks [Zou et al., 2025; Guo et al., 2021]. For unified and foundation region embedding, recent studies [Jenkins et al., 2019; Zhang et al., 2021] focus on learning general geospatial embeddings from a multimodal perspective. Nonetheless, efficiently integrating multimodal data from intricate geospaces, while maintaining generalizability, poses a significant challenge [Shang et al., 2020b; Zou et al., 2025; Rao et al., 2024; Shang et al., 2020a]. Specifically, conventional vision encoders (including multimodal large language models) trained on natural images find it challenging to grasp the geo-semantic details of satellite images, while the social impact from POIs in geospace is often overlooked. Additionally, popular modality fusion approaches in multimodal learning, like CLIPs [Hao et al., 2024; Yan et al., 2024], fall short in comprehending the cross-modal interactions between geo-entities within geospace. In this research, drawing inspiration from the success of semantic segmentation of satellite imagery [Zanaga et al., 2021] and heterogeneous graph methodologies, we propose an integrated geospatial embedding learning framework based on HG architecture that merges visual entities and POI attributes of geospace while considering mixed-order geospatial relationships with entities and regions.
Appendix B More Introduction of ESA WorldCover Dataset